# EndoChat: Grounded Multimodal Large Language Model for Endoscopic Surgery

### v1

@@ Results
EndoChat tailors interactions to varying complexity levels of surgical conversation.
@@
Single Phrase QA focuses on delivering concise and definitive answers, which is ideal for straightforward queries during surgical procedures. By employing the task-specific prompt: ”Answer the question with a single phrase.”, EndoChat provides succinct responses, including key aspects such as instrument types, counts, actions, or relative positions within the surgical scene. This capability relies on rapid analysis of visual content, ensuring efficiency without redundant elaboration. For instance, queries like “How many instruments are visible?” are answered directly, such as “three.” Empirical evaluations, shown in Table 1, demonstrate that EndoChat outperforms the stat-of-the-art MLLMs such as BiomedGPT and LLaVA-Med. Specifically, on the EndoVis-17 part, where other models fail to answer questions effectively (0% accuracy and F-score), EndoChat achieves a remarkable 55.51% accuracy, along with F-score (29.78), AP@50 (90.25), and mIoU (86.62). EndoChat also has superior performance on other parts of Surg-396K, and other public surgery datasets shown in Table 2 and 3, which demonstrates the robustness and effectiveness of the instruction-tuning process on our Surg-396K dataset.
Detailed Description caters to more complex scenarios where comprehensive understanding is necessary. This interaction type provides in-depth explanations grounded in the visual content of the surgical scene, making it essential for training scenarios and complex procedures. Answers generated by EndoChat offer detailed insights into tissues, instruments, motions, and other surgical elements, aiding in decision-making and contextual understanding, and also providing the reasoning to justify the answers. As shown in Table 4, EndoChat’s performance in generating detailed descriptions is rigorously assessed using GPT-4 Score, where it significantly outperforms all MLLMs across all parts of Surg-396K. This demonstrates EndoChat’s comprehensive advantage in generating detailed, context-aware, and high-quality descriptions.
@@
EndoChat enhances interactions through multiple modalities
@@
EndoChat demonstrates superior performance in Grounding QA, a crucial capability for real-time surgical guidance. By delivering responses through bounding boxes, Grounding QA enables EndoChat to ensure accurate spatial localization based on both the visual content and the posed questions. The task-specific prompt, ”Answer the question with just a bounding box.”, directs the model to focus on precise spatial information. As shown in Table 1 and 2, EndoChat outperforms other state-of-the-art models by a significant margin. Notably, EndoChat achieves the highest mIoU on both the EndoVis-18-VQLA and EndoVis-17-VQLA datasets, with scores of 86.89 and 86.62, respectively. The consistently superior performance of EndoChat over specialized models highlights its more effective integration of both visual and language-based reasoning, suggesting EndoChat is better suited for surgical navigation tasks.
@@
EndoChat is a surgeon-like interaction tool
@@
EndoChat advances comprehensive understanding of surgical scenerios
In order to comprehensively evaluate the capacity of EndoChat in addressing the challenges inherent in surgical environments, we conduct an in-context learning comparison between EndoChat and medical-specialized MLLMs in eight surgical scene understanding tasks on the CoPESD part of the Surg-396K dataset. The CoPESD part is chosen since it encompasses the full range of surgical understanding challenges. The tasks are formulated based on the dataset attributes illustrated in Figure 3 a, with each attribute defining one or two corresponding tasks [52]. These tasks reflect the essential components of surgical scenarios, such as instrument recognition, motion understanding, and issue detection. In-context learning is utilized for its advantage to dynamically adapt to new tasks and queries by leveraging task-specific prompts like "The answer must be one of the following words or phrases: ’Reach’, ’Rotate’, ’Grasp’, ’Lift’, ’Hold’, ’Stay idle’, ’Dissect’.". As summarized in Table 5, EndoChat consistently outperformed other medical-specialized MLLMs, showcasing its adaptability and precision in handling diverse understanding challenges.
@@
Table 5: Comparison experiments with zero-shot medical-specialized MLLMs in eight surgical scene understanding tasks on the CoPESD part of Surg-396K dataset. Eight types of tasks include: the number of instruments, the object location in the surgical scene (textual form), the current motion of the instrument, the direction of instruction motion, the identification of instruments, instrument detection, the recognition of issues and issue detection.
@@
Ablation study for the effectiveness of core modules in EndoChat
In the ablation study, we evaluate the effectiveness of our proposed modules in EndoChat: Mixed Visual Token Engine and Visual Contrast Hallucination Mitigation, utilizing the EndoVis-18 subset of the Surg-396K dataset. The results are shown in Table 6. For the evaluation of MVTE, we focus on three conversation types: Single Phrase, Grounding QA, and Detail Description since these are particularly sensitive to the quality of image feature extraction and perception. Specifically, for the Single Phrase, accuracy increases from 66.74% to 71.47% and F-score from 33.75% to 43.74%, demonstrating that MVTE significantly enhances the model’s ability to generate more accurate phrase-level descriptions. While the improvement in Grounding QA is modest, the increase in GPT-4 score (from 78.26 to 79.35) suggests that MVTE also contributes to refining the model’s reasoning capabilities, particularly for complex surgical scenarios. These results indicate that MVTE strengthens the model’s capacity to capture high-quality image features, thus improving performance in tasks requiring fine-grained image perception.
@@
Expert evaluation of EndoChat by endoscopists
To validate EndoChat’s potential in advancing surgical training and education, we conduct an expert evaluation involving 150 endoscopic surgery cases, evaluated by experienced endoscopists from Qilu Hospital. Each surgery case comprised a surgical image along with the corresponding five rounds of conversation. To ensure the evaluation’s comprehensiveness, the five rounds of conversation included a detailed description of the scenario, supplemented by four rounds of randomly selected Visual QA and Region Based QA, which encompass all attributes of the surgical data in Surg-396K. Additionally, the correct answers to these conversations are provided to assist the endoscopist in assessing EndoChat’s descriptions, analyses, and applicability in training scenarios.
@@
As shown in Figure 2 a, the scores ranged from strongly agree to strongly disagree, and 74.7% cases are evaluated as having correct descriptions provided by EndoChat, while 76.6% cases feature useful analysis that enhances the understanding of surgical scenes. Additionally, for 72% of the cases, endoscopists agreed that EndoChat could effectively assist trainees in surgical training, helping to refine procedural skills and improve educational outcomes. Finally, 69.3% of the cases reflected a willingness to incorporate EndoChat into surgical training, indicating its potential for real-world adoption. These findings highlight EndoChat’s role as a reliable tool and its effectiveness in advancing surgical training procedures and education in endoscopic surgery. Other than that, Figure 2 (b) illustrates the pairwise correlations among the evaluation standards. A strong positive correlation is observed between the correctness of answers and the willingness to use EndoChat, highlighting that its capacity to provide accurate and reliable information directly drives its potential for real-world implementation. Additionally, the training help exhibits a robust correlation with both the usefulness of analysis and correctness, indicating that practical and insightful analysis is pivotal for acceptance and education. These relationships emphasize that EndoChat’s ability to generate precise, contextually relevant outputs aligns with the endoscopists’ need for training support. Given the excellent performance demonstrated in previous experiments, EndoChat shows clear advantages in bridging the gap between AI innovation and endoscopic surgical practice through such correlations, further solidifying its value as a versatile and impactful tool.
Figure 2: Endoscopist evaluation of EndoChat in 150 cases. a Questionnaire-based evaluation of EndoChat conducted by endoscopists. The pie charts illustrate the distribution of cases that endoscopists express varying levels of agreement. b Correlation analysis of four evaluation standards.
Discussion
@@
Our analysis and comparisons include six MLLMs and over ten specialized models, demonstrating that our model achieved outstanding performance across various dialogue paradigms and surgical-specific scene understanding tasks. We further validate the effectiveness of our proposed visual feature learning approach and the visual contrast-based MLLM reasoning method through ablation experiments. Additionally, we invite experienced surgeons to evaluate their willingness to use EndoChat as an assistant during training or surgeries. In most evaluation cases, the surgeons provide positive feedback, further showcasing the clinical reliability, usability, and acceptability of our proposed EndoChat.
@@
Generally, there are two key factors in designing a precise and surgeon-friendly chatbot for dialogue and usage. The first is to ensure accuracy in downstream tasks for surgical scene understanding, such as instrument recognition and action identification. To enhance EndoChat’s performance in these downstream tasks, we consider single-phrase QA to be the most critical. This is because the answers in such dialogue data are often simple words or short phrases, making it easier for the model to link visual information with text annotations, thereby achieving higher accuracy in sub-tasks. The second is making the chatbot better suited to how surgeons use it, which is a primary focus of this study. Establishing different dialogue paradigms helps respond to questions from surgeons and trainees in various contexts, and also helps constrain the divergent dialogue tendencies typical of MLLMs. This allows the model to focus more on the questions posed by surgeons and provide relevant answers. In the future, we target to integrate EndoChat directly into surgical training or endoscopic surgery systems. By using monitors and voice-based dialogue systems, EndoChat could provide direct assistance to surgeons or trainees.
@@
Despite the impressive performance of our EndoChat on various surgical dialogue tasks, it still faces several limitations. First, although we possess a large surgical image database, the number of unique surgical cases included is relatively small. Such a large database facilitates tasks like action and instrument recognition; however, the limited number of cases may hinder the generalization ability of our model when applied to different surgical techniques. Expanding the database to include more surgical procedures and cases in the future could significantly enhance the generalizability and applicability of EndoChat across diverse surgical scenarios. In addition, MLLMs often rely on substantial computational power, which poses challenges for deployment in resource-constrained edge environments. Existing deployment approaches include developing lightweight MLLMs for deployment on mobile devices or personal computers, or hosting the model in the cloud and enabling communication with mobile/computer terminals. Finding ways to deploy MLLMs in clinical settings with limited computational resources will remain a significant challenge. Lastly, as more and more diverse data are introduced, issues concerning the privacy and ethical use of clinical data need to be carefully studied and reviewed to ensure compliance during their application.
@@
In conclusion, we present a flexible surgical understanding MLLM, EndoChat, designed to integrate various downstream tasks in surgical scene understanding and support different dialogue paradigms that may occur between surgeons and chatbots. Extensive experiments demonstrated the effectiveness of our approach, providing a more generalizable solution for understanding the surgical scene. Furthermore, we will open-source our model weights, training code, and data to promote the development of multimodal AI systems in the surgical domain. In the future, we will collaborate with surgeons and clinical systems to conduct more rigorous and extensive validations to ensure the safety, reliability, and usability of the dialogue model.
Methods
Surg-396K: Surgical Multimodal Instruction Dataset
The AI-assisted surgery field has experienced a notable expansion in the availability of public multimodal datasets, particularly visual-question-answer (VQA) pairs, as evidenced by works ranging from [6] to [10]. However, the availability of multimodal instruction data remains limited, primarily due to the time-intensive and less standardized processes involved in human crowd-sourcing. To promote the development of MLLMs tailored for surgical understanding, we propose Surg-396K, a surgical multimodal instruction dataset incorporating 41K images and 396K instruction-following annotation for endoscopic surgery. Following the data generation and aggregation process shown in Figure 3, we compile a total of 341k image-instruction pairs for training and 55k for testing. The subsequent subsections will detail the generation process of Surg-396K.
Preliminary for Constituent Datasets
In the construction of our Surg-396K dataset, we integrate three distinct datasets. Inspired by the achievements of recent MLLMs in text-annotation tasks [61], we utilize ChatGPT-4 to generate multimodal instruction-following data, resulting in five conversation types derived from EndoVis-VQLA [9] and CoPESD [49] datasets. The third dataset Cholec80-VQA [6], which lacks grounding information, is Surg-396K dataset.
Cholec80-VQA[6] is an innovative dataset generated from 40 video sequences of the Cholec80 dataset[64], encompassing a total of 21,591 frames. The images in Cholec80-VQA have a resolution of 854 × 480 pixels. Utilizing original tool-operation and phase annotations from the Cholec80 dataset, Cholec80-VQA proposes two types of question-answer pairs for each frame: Classification, which features 14 unique single-word answers; Sentence, which is presented in full sentence form. Due to the absence of grounding information and less content in the annotation of each image, we do not expand it with ChatGPT-4, but directly use Classification and Sentence as the conversation of the Single Phrase and Visual QA.
Attribute Extraction
To ensure that the annotations encompass comprehensive surgical information, we adopt a hierarchical framework for attribute analysis, from basic observation to dynamic operation and high-level perception, as shown in Figure 3 (a). At the Observation level, foundational attributes are defined, including Instrument Number (IN, count of visible instruments), Instrument Category (IC, classification of instrument types), and Target Issue (TI, anatomical targets of surgical focus). The Operation level focuses on dynamic behaviors and spatial characteristics, such as Object Position (OP, spatial mapping within a 3×3 grid), Instrument Motion (IM, functional roles inferred from motion), and Motion Direction (MD, trajectories across eight cardinal and diagonal directions). Finally, the Analysis level integrates these attributes to support semantic reasoning, encompassing tasks like Description and broader contextual reasoning. This structured design achieves a seamless integration of the dataset content that maximizes the extraction of surgical information from the original annotations. Furthermore, we develop attribute-specific QA templates tailored for the generation of instruction-tuning data. The distribution statistics for these attributes are shown in Figure 3 (c).
Category Sample
Attribute Instrument Number [IN] 2, 3, etc.
Instrument Category [IC] Prograsp forceps, ultrasound probe, etc.
Object Position [OP] Right-bottom, left-top, etc.
Instrument Motion [IM] Idle, lift, etc.
Target Issue [TI] Kidney, the mucosal flap, etc.
Motion Direction [MD] Upward, lower left, etc.
Category Response Style
Detail Description Describe the image in detail with [grounding].
Visual QA Question and answer without [grounding].
Grounding QA Answer the question with a [grounding].
Diverse Conversation Generation
The expression format of the instruction-tuning data obtained by information extraction through attributes is limited to human-designed templates, resulting in a homogeneous structure. In order to emulate the natural language expression, we further interpret the instruction-tuning data through ChatGPT-4 and generate five conversation types. The distribution statistics and style of these conversation types are shown in Figure 3 (d) and Table 7
Single Phrase is designed to equip EndoChat with the capability to deliver concise, definitive answers to each query, leveraging a rapid analysis of the surgical image. This type of conversation can be directly sourced from the instruction-tuning data. Additionally, we have enriched the diversity of the questions utilizing ChatGPT-4. To guide the model to provide the answer in a single word or phrase, we introduce the task-specific prompt “Answer the question with a single phrase.” at the end of the questions, which can be represented by:
[…]
Detailed Description provides comprehensive, grounded responses to queries that delve into the intricate details of the visual scene. Answers in this type are entirely generated by ChatGPT-4 with guided prompts that take full advantage of instruction-tuning data. Therefore, we ensure the response covers all attributes in the corresponding images. The list of questions has also been diversified using ChatGPT-4. This type can augment EndoChat’s proficiency in articulating complex visual information as if it is observing the scene in real time.
Visual QA emphasizes straightforward question-answer pairs that provide general insights into the surgical scene without specific grounding. This mode differs from Single Phrase by allowing more contextual information in the responses while still maintaining conciseness. Therefore, the generation process for Visual QA entails an additional step compared to Single Phrase. Specifically, ChatGPT-4 is utilized to elaborate the single-word or single-phrase response into a complete sentence while incorporating the content related to description and reasoning attributes from the instruction-tuning data.
Region Based QA incorporates grounding information within the question to guide the model’s attention to a specific region of the image. This conversation type facilitates targeted analysis of visual content, pinpointing the precise location of surgical instruments or the condition of specific tissues compared to Visual QA. For text expression, we insert the bounding box of the target after its text, e.g., "kidney [$x_{1},y_{1},x_{2},y_{2}$]". $x_{1}$ and $y_{1}$ denote the coordinates of the top-left corner of the bounding box, while $x_{2}$ and $y_{2}$ specify the bottom-right corner. Each coordinate value is normalized to the interval $[0,1]$.
Grounding QA delivers responses solely through bounding boxes, training EndoChat to provide accurate spatial answers based on the interplay between the visual content and the posed questions. We also introduce the task-specific prompt “Answer the question with just a bounding box.” at the end of the questions to guide the model to provide the answer in a bounding box. This conversation type has the same format as Single Phrase QA in Equation 1.
Data Cleaning
Following the generation of diverse conversations, a data cleaning process is implemented to ensure the integrity and reliability of the training data. Given the large scale of the Surg-396K, we conduct the sampling-based inspection with a ratio of 1/5. Specifically, we manually review the sampled text to evaluate its information completeness, relevance, and clarity. Information completeness refers to whether the text includes all essential content, such as operations and tools. We assign conversation type labels to each QA pair and assess relevance by verifying whether the content aligns with its assigned labels. Since the QA pairs are semantically enriched using ChatGPT-4, we also inspect whether the enriched text is semantically clear and accurate. During the data cleaning process, we document frequently occurring issues and apply modifications to non-sampled content accordingly. This procedure ensures that retained and revised annotations can provide meaningful information for MLLM training.
EndoChat
Architecture of EndoChat
Visual grounding conversations in endoscopic surgery involve the intricate interaction between visual and linguistic modalities, requiring a comprehensive understanding of knowledge about distinct objects or regions within the endoscopic surgery scenes. Therefore, we propose EndoChat, a novel large vision-language model capable of engaging in visual grounding conversations concerning endoscopic surgery. Given an input image, the mixed visual encoder first extract source tokens, denoted as $X_{d}\in\mathbb{R}^{N\times D\times L_{1}}$ and $X_{o}\in\mathbb{R}^{N\times D\times L_{2}}$, where $N$ represents the number of frames, $D$ denotes the hidden dimension, and $L_{1}$ and $L_{2}$ correspond to the sequence lengths of the respective token sets. Subsequently, the extracted source tokens are processed by our proposed mixed visual token engine, which augments the representation by generating additional visual tokens. The resulting enhanced image tokens are represented as $X^{\prime}\in\mathbb{R}^{N\times(D+m)\times(L_{1}+L_{2})}$, where $m$ denotes the number of newly generated tokens. These enriched visual tokens are then aligned with language tokens and input to the language model to produce the final response. Moreover, we introduce a visual contrast mechanism to mitigate object hallucinations, further enhancing the model’s visual understanding and ensuring consistency between visual inputs and language outputs in complex endoscopic surgery scenarios.
Mixed Visual Token Engine
In our EndoChat, we follow SPHINX that mixes visual embedding to more scales from high-resolution sub-images, thereby enhancing the encoding of high-resolution images. For input images with high resolution, we implement two parallel pathways to generate five corresponding images at resolutions of 224×224 and 512×512, respectively. Then, these images are fed into a mixed visual encoder, which consists of DINOv2 [65] and OpenCLIP [66], resulting in outputs $X_{d}$ and $X_{o}$. For MLLMs, visual encoders typically summarize the visual embeddings after encoding image tokens by extracting an aggregated representation through operations like multi-layer perceptron (MLP). Although this direct representation is computationally efficient, it struggles to capture multi-scale information and often overlooks crucial spatial relationships between different positions or regions. Thus, it may confuse the LLM and underutilize its capabilities. To address these limitations, we introduce the Mixed Visual Token Engine (MVTE). MVTE dynamically generates global visual tokens based on the source token produced by the mixed visual encoder, which seamlessly integrates and maximizes the informational utility of multi-scale visual tokens.
$X^{\prime}=\text{Softmax}(\text{MLP}(X))\cdot X\oplus X$ (2)
$\hat{X}=\text{Proj}(X_{o}^{\prime}\oplus X_{d}^{\prime})$ (3)
Hallucination Mitigation through Visual Contrast
Implementation Details
We adopt the open-source LLaMA2-13B [50] large language model in the SPHINX architecture as our EndoChat’s foundational component. LLaMA2-13B serves as a unified interface for diverse vision-language tasks. To ensure the model’s responses are aligned and contextually effective, task-specific prompts are appended to input data, which helps guide the LLM’s responses. For LLM fine-tuning, we employ the Low-Rank Adaptation (LoRA) [70] technique that introduces two smaller matrices as a low-rank approximation of the large, original matrix. We optimize parameters of low-rank matrices instead of all parameters in the pre-trained LLaMA-2. This adaptation method reduces training time and computational overhead. At the same time, it preserves the model’s broader knowledge of generic object categories and spatial landmarks, thereby enhancing its vision-language reasoning capabilities in the endoscopic surgery domain. To train the LLM and the mixed visual token engine, we utilize an input resolution of 1024×1024. The training process is conducted on the Surg-396K dataset for a single epoch, employing four NVIDIA A800 GPUs. The initial learning rate is set to 2e-5 with a batch size of 16, requiring about 20 hours to complete.
Data availability
The Surg-396K dataset is publicly available at https://github.com/gkw0010/EndoChat.
Code availability
The codes of EndoChat can be downloaded at https://github.com/gkw0010/EndoChat.
Acknowledgements

### v2

@@ 4.1 Implementation Details
4.2 EndoChat tailors interactions to varying complexity levels of surgical conversation.
@@
Single Phrase QA focuses on delivering concise and definitive answers, which is ideal for straightforward queries during surgical procedures. By employing the task-specific prompt: ”Answer the question with a single phrase.”, EndoChat provides succinct responses, including key aspects such as instrument types, counts, actions, or relative positions within the surgical scene. This capability relies on rapid analysis of visual content, ensuring efficiency without redundant elaboration. For instance, queries like “How many instruments are visible?” are answered directly, such as “three.” Empirical evaluations, shown in Table 3, demonstrate that EndoChat outperforms the state-of-the-art MLLMs such as BiomedGPT and LLaVA-Med. Specifically, on the EndoVis-17 part, where other models fail to answer questions effectively (0% accuracy and F-score), EndoChat achieves a remarkable 55.51% accuracy, along with F-score (29.78), AP@50 (90.25), and mIoU (86.62). It also has superior performance on other parts of Surg-396K, and other public surgery datasets shown in Table 4 and 5, which demonstrates the robustness and effectiveness of the instruction-tuning process on our Surg-396K dataset.
Detailed Description caters to more complex scenarios where comprehensive understanding is necessary. This interaction type provides in-depth explanations grounded in the visual content of the surgical scene, making it essential for training scenarios and complex procedures. Answers generated by EndoChat offer detailed insights into tissues, instruments, motions, and other surgical elements, aiding in decision-making and contextual understanding, and also providing the reasoning to justify the answers. As shown in Table 6, EndoChat’s performance in generating detailed descriptions is rigorously assessed using GPT-4 Score, where it significantly outperforms all MLLMs across all parts of Surg-396K. This demonstrates EndoChat’s advantage in generating detailed, context-aware, and high-quality descriptions.
@@
4.3 EndoChat enhances interactions through multiple modalities
@@
EndoChat demonstrates superior performance in Grounding QA, a crucial capability for real-time surgical guidance. By delivering responses through bounding boxes, Grounding QA enables EndoChat to ensure accurate spatial localization based on both the visual content and the posed questions. The task-specific prompt, ”Answer the question with just a bounding box.”, directs the model to focus on precise spatial information. As shown in Table 3 and 4, EndoChat outperforms other state-of-the-art models by a significant margin. Notably, EndoChat achieves the highest mIoU on both the EndoVis-18-VQLA and EndoVis-17-VQLA datasets, with scores of 86.89 and 86.62, respectively. The consistently superior performance of EndoChat over specialized models highlights its more effective integration of both visual and language-based reasoning, suggesting EndoChat is better suited for surgical navigation tasks.
@@
4.4 EndoChat is a surgeon-like interaction tool
@@
4.5 EndoChat advances comprehensive understanding of surgical scenerios
In order to comprehensively evaluate the capacity of EndoChat in addressing the challenges inherent in surgical environments, we conduct an in-context learning comparison between EndoChat and medical-specialized MLLMs in seven surgical scene understanding tasks on the CoPESD part of the Surg-396K dataset. The CoPESD part is chosen since it encompasses the full range of surgical understanding challenges. The tasks are formulated based on the dataset attributes illustrated in Figure 2 a, with each attribute defining one or two corresponding tasks [42]. These tasks reflect the essential components of surgical scenarios, such as instrument recognition, motion understanding, and issue detection. In-context learning is utilized for its advantage to dynamically adapt to new tasks and queries by leveraging task-specific prompts like "The answer must be one of the following words or phrases: ’Reach’, ’Rotate’, ’Grasp’, ’Lift’, ’Hold’, ’Stay idle’, ’Dissect’.". As summarized in Table 7, EndoChat consistently outperformed other medical-specialized MLLMs, showcasing its adaptability and precision in handling diverse understanding challenges.
@@ 4.5 EndoChat advances comprehensive understanding of surgical scenerios
Table 7: Comparison experiments with zero-shot medical-specialized MLLMs in various surgical scene understanding tasks (excluding the Description task which has been shown in Table 6) on the CoPESD part of Surg-396K dataset. Six types of tasks include: the number of instruments, the object location in the surgical scene (textual form), the current motion of the instrument, the direction of instruction motion, the identification of instruments, instrument detection, the recognition of issues and issue detection.
@@
4.6 Ablation study for the effectiveness of core modules in EndoChat
In the ablation study, we evaluate the effectiveness of our proposed modules in EndoChat: Mixed Visual Token Engine and Visual Contrast Hallucination Mitigation, utilizing the EndoVis-18 subset of the Surg-396K dataset. The results are shown in Table 8. For the evaluation of MVTE, we focus on three conversation types: Single Phrase, Grounding QA, and Detail Description, since these are particularly sensitive to the quality of image feature extraction and perception. Specifically, for the Single Phrase, accuracy increases from 66.74% to 71.47% and F-score from 33.75% to 43.74%, demonstrating that MVTE significantly enhances the model’s ability to generate more accurate phrase-level descriptions. While the improvement in Grounding QA is modest, the increase in GPT-4 score (from 78.26 to 79.35) suggests that MVTE contributes to refining the model’s reasoning capabilities, particularly for complex surgical scenarios. These results indicate that MVTE strengthens the model’s capacity to capture high-quality image features, thus improving performance in tasks requiring fine-grained image perception.
@@ 4.6 Ablation study for the effectiveness of core modules in EndoChat
4.7 Expert evaluation of EndoChat by endoscopists
To validate EndoChat’s potential in advancing surgical training and education, we conduct an expert evaluation involving 150 endoscopic surgery cases, evaluated by experienced endoscopists from Qilu Hospital. Each surgery case comprised a surgical image along with the corresponding five rounds of conversation. To ensure the evaluation’s comprehensiveness, the five rounds of conversation included a detailed description of the scenario, supplemented by four rounds of randomly selected Visual QA and Region Based QA, which encompass all attributes of the surgical data in Surg-396K. Additionally, the ground-truth to these conversations are provided to assist the endoscopist in assessing EndoChat’s descriptive accuracy, analytical depth, and applicability in training scenarios.
@@ 4.7 Expert evaluation of EndoChat by endoscopists
As shown in Figure 4 a, the scores ranged from strongly agree to strongly disagree, and 74.7% cases are evaluated as having correct descriptions provided by EndoChat, while 76.6% cases feature useful analysis that enhances the understanding of surgical scenes. Additionally, for 72% of the cases, endoscopists agreed that EndoChat could effectively assist trainees in surgical training, helping to refine procedural skills and improve educational outcomes. Finally, 69.3% of the cases reflected a willingness to incorporate EndoChat into surgical training, indicating its potential for real-world adoption. These findings highlight EndoChat’s role as a reliable tool and its effectiveness in advancing surgical training procedures and education in endoscopic surgery. Other than that, Figure 4 (b) illustrates the pairwise correlations among the evaluation standards. A strong positive correlation is observed between the correctness of answers and the willingness to use EndoChat, highlighting that its capacity to provide accurate and reliable information directly drives its potential for real-world implementation. Additionally, the training helps exhibit a robust correlation with both the usefulness of analysis and correctness, indicating that practical and insightful analysis is pivotal for acceptance and education. These relationships emphasize that EndoChat’s ability to generate precise, contextually relevant outputs aligns with the endoscopists’ need for training support. Given the excellent performance demonstrated in previous experiments, EndoChat shows clear advantages in bridging the gap between AI innovation and endoscopic surgical practice through such correlations, further solidifying its value as a versatile and impactful tool.
Figure 4: Endoscopist evaluation of EndoChat in 150 cases. a Questionnaire-based evaluation of EndoChat conducted by endoscopists. The pie charts illustrate the distribution of cases in which endoscopists express varying levels of agreement. b Correlation analysis of four evaluation standards.
5 Discussion
@@
Our analysis and comparisons include six MLLMs and over ten specialized models, demonstrating that our model achieved outstanding performance across various dialogue paradigms and surgical-specific scene understanding tasks. We further validate the effectiveness of our proposed visual feature learning approach and the visual contrast-based MLLM reasoning method through ablation experiments. Additionally, we invite experienced endoscopists to evaluate their willingness to use EndoChat as an assistant during training or surgeries. In most evaluation cases, the surgeons provide positive feedback, further showcasing the clinical reliability, usability, and acceptability of our proposed EndoChat. There are demonstrations of qualitative comparison in supplementary material.
@@
Generally, there are two key factors in designing a precise and surgeon-friendly chatbot for dialogue and usage. The first is to ensure accuracy in downstream tasks for surgical scene understanding, such as instrument recognition and action identification [60]. To enhance EndoChat’s performance in these downstream tasks, we consider single-phrase QA to be the most critical. This is primarily due to the answers in such dialogue data are often simple words or short phrases, making it easier for the model to link visual information with text annotations, thereby achieving higher accuracy in sub-tasks. The second is making the chatbot better suited to how surgeons use it, which is a main focus of this study. Establishing different dialogue paradigms helps respond to questions from surgeons and trainees in various contexts, and also helps constrain the divergent dialogue tendencies typical of MLLMs. This allows the model to focus more on the questions posed by surgeons and provide relevant answers.
@@
Despite the impressive performance of our EndoChat on various surgical dialogue tasks, it still faces several limitations. First, although we possess a large surgical image database, the number of unique surgical cases included is relatively small. Such a large database facilitates tasks like action and instrument recognition; however, the limited number of cases may hinder the generalization ability of our model when applied to different surgical techniques [78, 25]. Expanding the database to include more surgical procedures and cases in the future could significantly enhance the generalizability and applicability of EndoChat across diverse surgical scenarios. In addition, MLLMs often rely on substantial computational power, which poses challenges for deployment in resource-constrained edge environments. Existing deployment approaches include developing lightweight MLLMs for deployment on mobile devices or personal computers, or hosting the model in the cloud and enabling communication with mobile/computer terminals [76, 84]. Finding ways to deploy MLLMs in clinical settings with limited computational resources will remain a significant challenge. Lastly, as more and more diverse data are introduced, issues concerning the privacy and ethical use of clinical data need to be carefully studied and reviewed to ensure compliance during their application.
6 Conclusion
In this paper, we introduce Surg-396K, a comprehensive surgical multimodal dataset that includes 396K image-instruction pairs across multiple conversation paradigms. Based on Surg-396K, we present a flexible surgical understanding MLLM, EndoChat, designed to integrate various downstream tasks in surgical scene understanding and support different dialogue paradigms that may occur between surgeons and chatbots. Endochat integrates the Mixed Visual Token Engine (MVTE) and the visual contrast-based hallucination mitigation strategy, allowing it to effectively capture high-quality visual features and reduce inaccuracies in generated responses. Extensive experiments demonstrated the effectiveness of our approach, providing a more generalizable solution for understanding the surgical scene. Furthermore, the positive feedback from expert evaluations underscores the model’s practical applicability in real-world surgical environments.
@@ 6 Conclusion
Moving forward, we will open-source our model weights, training code, and data to promote the development of multimodal AI systems in the surgical domain. In the future, we will collaborate with surgeons and clinical systems to conduct more rigorous and extensive validations to ensure the safety, reliability, and usability of the dialogue model. We target to integrate EndoChat directly into surgical training or endoscopic surgery systems. By using monitors and voice-based dialogue systems, EndoChat could provide direct assistance to surgeons or trainees.
7 Acknowledgements